OpenStudy (anonymous):

Solve for x: Log base 10 (x) / Log base 10 (2) = 4

OpenStudy (anonymous):

x=16

OpenStudy (anonymous):

Does 4 become log base 10 (10000)?

OpenStudy (anonymous):

How did you get 16?

OpenStudy (anonymous):

2^4 = 16

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{\log _{10} x }{\log _{10} 2} =4 } \]\[\huge\color{blue}{ \log _{2} x =4 } \]\[\huge\color{blue}{ \log _{2} x =4 } ~~~~~~->~~~~~x=2^4\]

OpenStudy (anonymous):

How did it change from base 10 to base 2?

OpenStudy (anonymous):

What is the rule?

OpenStudy (solomonzelman):

I used a formula. \[\huge\color{blue}{ \frac{Log_{10}b}{Log_{10}a} = Log _{a} b } \]

OpenStudy (anonymous):

Thanks for explaining it to me. I get it now

OpenStudy (solomonzelman):

yw